Karina has a piece of wire that is 16 inches long. Use this fact to answer the following questions. Round your answers to the nearest tenth.
a. What will be the area inside a semicircle made from this wire?
c. Suppose that Karina cuts the wire into 2 pieces to make 2 squares such that the area of 1 square is twice the area of the other. What is the side length of the larger square?
thanks!

Answer by stanbon(75887)

Karina has a piece of wire that is 16 inches long. Use this fact to answer the following questions. Round your answers to the nearest tenth.
a. What will be the area inside a semicircle made from this wire?
Perimeter = (pi)d = 16 inches
d = 16/(pi)
radius = 8/(pi)
----------------------------------
Area = (pi)[8/(pi)]^2
Area = 64/(pi) sq. inches
=================================================
c. Suppose that Karina cuts the wire into 2 pieces to make 2 squares such that the area of 1 square is twice the area of the other. What is the side length of the larger square?
Let one piece be "x":
Other piece is "16-x"
-----
Areas:
1st square
side = x/4
Area = (x/4)^2
-----
2nd square:
side = (16-x)/4
Area = [(16-x)/4]^2
------
Equation:
Area #1 = 2 Area #2
(x/4)^2 = 2[(16-x)/4]^2
-------------------
Multiply both sides by 16:
x^2 = 2(16-x)^2
x^2 = 2(256 - 32x + x^2)
x^2 = 512 - 64x + 2x^2
---
x^2 - 64x +512 = 0
I graphed it to get x = 9.37 and x = 54.63
===============================================
Cheers,
Stan H.
===========================




thanks!